# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 5. The error term: Initial Transformations
Now we begin analyzing the sum $B_{w}(X,r)$. Interchanging the order of summation and applying Lemma 2.2 to the sum over $c_{1}$,
@@
The lemma below shows that $\widehat{F_{a_{1},c_{1}}}\left(\frac{n}{a_{1}}\right)$ is negligibly small if $n\gg\frac{X^{1+\varepsilon}}{lH}$ and $\frac{1}{l}\widehat{F_{la_{1},x}}\left(\frac{m}{la_{1}}\right)$ is negligibly small if $m\gg\frac{X^{1+\varepsilon}}{H}$.
@@
for every integer $k\geq 1$, where $F_{a,c}(x)$ is defined in (3.11). In particular, by taking $k=1$, we have the bound
@@
Note that $w^{(k)}\left(\frac{cx}{a}\right)=0$ unless $\frac{cx}{a}\in(1-H,1)\cup(X,X+H)$, and so $w^{(k)}\left(\frac{cx}{a}\right)$ vanishes outside two intervals of length $O\left(\left|\frac{a}{c}\right|H\right)$ each and by (3.7), we obtain the bound $O({\left|\frac{c}{a}\right|}^{k-1}y^{-k}H^{1-k})$ for the term corresponding to $j=0$. Now we consider the other summands. Since $w^{(j)}(x)$ is non-zero only in the intervals $(1-H,1)$ and $(X,X+H)$ for every $j\geq 1$, all these terms together contribute only $O(y^{-k}H^{1-k})$, again by (3.7). Finally, by applying the bound (5.3) in one of the integrals and by trivially bounding the other integral using the support of the function $w$, we obtain the bound (5.4). ∎
Now we isolate the $m=0$ term in $B_{w}(X,r)$ and write
$B_{w}(X,r)=B_{0,w}(X)+B_{1,w}(X),$ (5.5)
@@
5.1. Estimation of $B_{0,w}(X)$
Note that the Kloosterman sum appearing in $B_{0,w}(X)$ is nothing but the Ramanujan sum $r_{a_{1}}(nr_{1})$, and by the bound $|r_{q}(n)|\leq(q,n)$ (see [IK04, Eq. (3.5)]), we easily obtain the following:
Proposition 5.2.
We have,
$B_{0,w}(X)\ll X^{1+\varepsilon}.$ (5.6)
@@ 5.2. Analysis of $B_{1,w}(X)$: initial steps
$\mathcal{E}(Y)$ $:=\sum_{l|r}\sum_{0\neq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{0\neq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}a_{1}>0\\ la_{1}\in(Y,Y+H]\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}^{2}}\widehat{G}(n/a_{1},m/a_{1}).$
@@
This follows easily from (5.4) and the Weil bound (Lemma 2.3). ∎
@@
In view of the above proposition, whenever we have a sum over $a_{1}$, we can disregard the portion $\sum_{la_{1}>X}$ by adding the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$. Thus, from now onwards, whenever we encounter a sum of the above form, we may assume that the variable $a_{1}$ lies in the interval $[1,X]$ but we may not display the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ on every occasion since this term is already present in the error term in the final result, i.e., Prop. 3.3.
@@
$\Delta(\alpha,\beta,\gamma,l,U)$ $:=\sum_{\begin{subarray}{c}0\neq n\ll\frac{X^{1+\varepsilon}}{lH}\end{subarray}}\frac{1}{n^{\gamma}}\sum_{\begin{subarray}{c}0\neq m\ll\frac{X^{1+\varepsilon}}{H}\end{subarray}}\frac{1}{m^{\beta}}\sum_{\begin{subarray}{c}1\leq la_{1}\leq U\end{subarray}}\frac{w(la_{1})\left|S(nr_{1},-m,a_{1})\right|}{a_{1}^{\alpha}}$
@@ 5.3. Decomposition of $\widehat{G}(n/a_{1},m/a_{1})$
where we use Prop. 5.5 to estimate the contribution of the error term. From $I_{4}$ and $I_{5}$ we extract stationary phases for further analysis. The other terms can be handled by the Weil bound and they contribute only to the error term. Thus we obtain the following intermediate proposition, where we have used the notation
$\mathop{\sum_{l}\sum_{m}\sum_{n}}:=\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}.$ (5.9)
Proposition 5.6.
We have,
$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}$
$+\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right).$
say.
We proceed to prove the proposition below and the proof ends in §7.8.
@@ 6. Contributions of $I_{1},I_{2},I_{3},$ and $I_{6}$
It is enough to estimate the integrals $I_{1},I_{2},I_{3},$ and $I_{6}$ since their contributions can be estimated easily by the Weil bound or using Prop. 5.5.
We first prove (iii) and consider only $I_{2}$ since $I_{6}$ can be handled in a similar manner. To bound $I_{2}$, we first interchange the $x$ and $y$ integrals and then use integration by parts in the $x$-integral, obtaining
$I_{2}$ $=-\frac{1}{2\pi i}\frac{a_{1}}{n}\int\limits_{\frac{X}{l}}^{\frac{X+H}{l}}w(ly)e\left(\frac{-my}{a_{1}}\right)\int\limits_{x\leq la_{1}}\left(w^{\prime}(x)w\left(\frac{xy}{a_{1}}\right)+\frac{y}{a_{1}}w(x)w^{\prime}\left(\frac{xy}{a_{1}}\right)\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y+O\left(\frac{a_{1}H}{nl}\right)$
$\ll\frac{a_{1}H}{nl},$
@@
${I_{3}}^{\ast}$ $:=\int\limits_{0}^{a_{1}(H-1)}w(-ly)e\left(\frac{my}{a_{1}}\right)\int\limits_{1}^{\min\{la_{1},a_{1}(H-1)/y)\}}w\left(\frac{-xy}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y$
@@
For ${I_{3,2}}^{\ast}$, we note that from the support of the function $w$ (see (3.4)), $w(-ly)=0$ if $y>(H-1)/l$. Hence,
@@
The contribution of the $O$-terms is $O(X^{3/2+\varepsilon})$ by Prop. 5.5. We now analyze the exponential integral
@@
Let us assume further $m,n>0$ since the analysis in the other case is similar and we get the same bounds. We consider two cases.
Case I: $x_{0}\in[2,2la_{1}/3]$. Here we split the interval $[1,la_{1}]$ in three parts as follows:
@@ 7.1. Contributions of $I_{4}$
By Prop. 5.5, the contributions of the $O$-terms are absorbed into the error term. Define
$I_{m,n,a_{1}}:=\int_{la_{1}}^{X}e(h(x))\,\textnormal{d}x,\quad\textnormal{ where }\quad h(x):=-\frac{nx}{a_{1}}-\frac{mX}{x}.$
@@
If $mn<0$, then there is no stationary point and $|h^{\prime}(x)|\geq\frac{n}{a_{1}}$ throughout the interval and in this case we use the first derivative bound directly to the integral $I_{m,n,a_{1}}$. This gives $I_{m,n,a_{1}}\ll\frac{a_{1}}{n}$ and the contribution of this is sufficiently small by Prop. 5.5. The case left to consider is $mn>0$. In this case the stationary point is given by
@@
We will assume that $m,n>0$ in our analysis since the other case is identical. Now, if we give an upper bound for exponential integrals and then sum over the three variables $m,n,$ and $a_{1}$ as in the case of $I_{1}$ then the error term we get is not good enough. We, therefore, refine our analysis by localizing the stationary point $x_{0}$ and extracting the stationary phases wherever they occur and estimating the ensuing error terms carefully. This leads to the following:
@@
Before we begin the proof, let us first observe that from (7.2), we record the following crude range for $x_{0}$: $x_{0}\ll X\sqrt{X/H}\ll X^{5/4}$.
@@ 7.2. Localizing the stationary point and decomposition of $B_{1,w}(X,I_{4})$
Let us define, for a fixed $a_{1}$,
$k=k(a_{1})$ $:=\max\left\{j\geq 0:\frac{X}{2^{j}}\geq la_{1}+1\right\}=\left[\frac{\log(X/(la_{1}+1))}{\log 2}\right].$ (7.3)
@@
for $\text{for }0\leq j\leq k=k(a_{1})\leq J(X),\text{ we have }X/2^{j}\geq la_{1},\text{ where }J(X):=\left[\frac{\log(X/2)}{\log 2}\right].$ (7.4)
We make a dyadic subdivision of the range of $x_{0}$, i.e., the interval $[1,X^{5/4}]$ in the following manner:
$\mathcal{I}_{0}:=\left[X+\delta,X^{\frac{5}{4}}\right],\mathcal{I}_{j}:=\left[\frac{X}{2^{j}}+\delta,\frac{X}{2^{j-1}}-\delta\right]\text{ for }1\leq j\leq k,\text{ and }\mathcal{I}_{k+1}:=[la_{1}+\delta,\frac{X}{2^{k}}-\delta],\mathcal{I}_{k+2}:=[1,la_{1}-\delta].$ (7.5)
@@
Here $\delta>0$ is to be chosen later (we will see that $\delta=1/100$ suffices). Clearly, we can write the interval $[1,X^{5/4}]$ as a disjoint union of the $\mathcal{I}_{j}$’s and $\mathcal{I}^{\prime}_{j}$s, and thus we express $B_{1,w}(X,I_{4})$ as
$B_{1,w}(X,I_{4})$ $=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{0\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$ (7.7)
$+\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{1\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{x_{0}\in\mathcal{I}_{j}^{\prime}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$
@@
say.
7.3. Subdividing the range of integration and decomposition of the sum $\mathcal{S}$
We first subdivide the range of the integral $[la_{1},X]$. Let
$\mathcal{F}_{j}:=\left[\frac{X}{2^{j}},\frac{X}{2^{j-1}}\right],\text{ for }1\leq j\leq k\text{ and }\mathcal{F}_{k+1}:;=\left[la_{1},\frac{X}{2^{k}}\right].$ (7.8)
Clearly, $[la_{1},X]=\bigcup\limits_{1\leq j\leq k+1}\mathcal{F}_{j},$ where an interval $\mathcal{F}_{j}$ can intersect its neighbour only at the boundary point. Let us define

### v2

@@ 5. The error term: Initial Transformations
Our goal now is to obtain the following intermediate proposition, where we have used the notation
$\mathop{\sum_{l,m,n}}:=\sum_{l|r}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}.$ (5.1)
Proposition 5.1.
We have,
$B_{w}(X,r)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{3}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnX}{a_{1}}}-\frac{1}{8}\right)$
$-e\left(2\sqrt{\frac{mnX}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}$
$+\frac{1}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}\sum_{\begin{subarray}{c}0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}\left(\frac{a_{1}}{mn}\right)^{\frac{1}{4}}\bigg{\{}e\left(-2\sqrt{\frac{mnu}{a_{1}}}-\frac{1}{8}\right)$
$+e\left(2\sqrt{\frac{mnu}{a_{1}}}+\frac{1}{8}\right)\bigg{\}}+O\left(X^{\varepsilon}\left(X^{\frac{3}{2}}+X^{\frac{5}{4}}\sqrt{\frac{X}{H}}+\frac{X^{2}}{H}+\left(\frac{X}{H}\right)^{3}+\frac{X^{\frac{5}{2}}}{H^{2}}\right)\right)$
say.
We proceed to prove the proposition and the proof ends in §7.8.
Interchanging the order of summation and applying Poisson summation to the sum over $c_{1}$ after subdividing into residue classes modulo $a_{1}$, we have
@@
The decay of $\widehat{G}$ limits the effective ranges of the variables $m$ and $n$ to $\frac{X^{1+\varepsilon}}{H}$ and $\frac{X^{1+\varepsilon}}{lH}$, respectively, as shown in the lemma below. First recall the definition in (3.10).
@@
for every integer $k\geq 1$. In particular, by taking $k=1$, we have the bound
@@
Note that $w^{(k)}\left(\frac{cx}{a}\right)=0$ unless $\frac{cx}{a}\in(1-H,1)\cup(X,X+H)$, and so $w^{(k)}\left(\frac{cx}{a}\right)$ vanishes outside two intervals of length $O\left(\left|\frac{a}{c}\right|H\right)$ each and by (3.7), we obtain the bound $O({\left|\frac{c}{a}\right|}^{k-1}y^{-k}H^{1-k})$ for the term corresponding to $j=0$. Now we consider the other summands. Since $w^{(j)}(x)$ is non-zero only in the intervals (1H,1)(1-H,1) and $(X,X+H)$ for every $j\geq 1$, all these terms together contribute only $O(y^{-k}H^{1-k})$, again by (3.7). Finally, by applying the bound (5.4) in one of the integrals and by trivially bounding the other integral using the support of the function $w$, we obtain the bound (5.5). ∎
Now we isolate the $m=0$ term in $B_{w}(X,r)$. As the Kloosterman sum becomes a Ramanujan sum which satisfies the bound $|r_{q}(n)|\leq(q,n)$ (see [IK04, Eq. (3.5)]), we easily obtain the following:
$B_{w}(X,r)=B_{1,w}(X)+O(X^{1+\varepsilon}),$ (5.6)
@@ 5.1. Analysis of $B_{1,w}(X)$: initial steps
$\mathcal{E}(Y)$ $=\sum_{l|r}\sum_{0\neq n\ll\frac{X^{1+\varepsilon}}{lH}}\sum_{0\neq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{\begin{subarray}{c}a_{1}>0\\ la_{1}\in(Y,Y+H]\end{subarray}}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}^{2}}\widehat{G}(n/a_{1},m/a_{1}).$
@@
This follows easily from (5.5) and the Weil bound for Kloosterman sums. ∎
@@
In view of the above proposition, wherever we have a sum over $a_{1}$, we can disregard the part $\sum_{la_{1}>X}$ by adding the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$. Thus, from now onwards, whenever we encounter a sum of the above form, we may assume that the variable $a_{1}$ lies in the interval $[1,X]$ but we may not display the term $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ on every occasion since this term is already present in the error term in the final result, i.e., Prop. 3.2.
@@
$\Delta(\alpha,\beta,\gamma,l,U)$ $=\sum_{\begin{subarray}{c}0\neq n\ll\frac{X^{1+\varepsilon}}{lH}\end{subarray}}\frac{1}{n^{\gamma}}\sum_{\begin{subarray}{c}0\neq m\ll\frac{X^{1+\varepsilon}}{H}\end{subarray}}\frac{1}{m^{\beta}}\sum_{\begin{subarray}{c}1\leq la_{1}\leq U\end{subarray}}\frac{w(la_{1})\left|S(nr_{1},-m,a_{1})\right|}{a_{1}^{\alpha}}$
@@ 5.2. Decomposition of $\widehat{G}(n/a_{1},m/a_{1})$
where we use Prop. 5.4 to estimate the contribution of the error term. From $I_{4}$ and $I_{5}$ we extract stationary phases for further analysis. The other terms can be handled by the Weil bound and they contribute only to the error term.
@@ 6. Contributions of $I_{1},I_{2},I_{3},$ and $I_{6}$
It is enough to bound the integrals as the outside sums are estimated trivially after applying the Weil bound (see Prop. 5.4). We first prove (iii) and consider only $I_{2}$ since $I_{6}$ can be handled in a similar manner. Interchanging the order of the two integrals appearing in $I_{2}$ and using integration by parts in the $x$-integral, we get
$I_{2}\ll\frac{a_{1}H}{nl},$
@@
${I_{3}}^{\ast}$ $=\int\limits_{0}^{a_{1}(H-1)}w(-ly)e\left(\frac{my}{a_{1}}\right)\int\limits_{1}^{\min\{la_{1},a_{1}(H-1)/y)\}}w\left(\frac{-xy}{a_{1}}\right)e\left(\frac{-nx}{a_{1}}\right)\textnormal{d}x\textnormal{d}y$
@@
For $I_{3,2}$$^{\ast}$, we note that from the support of the function $w$ (see (3.4)), w(ly)=0w(-ly)=0 if y>(H1)/ly>(H-1)/l. Hence,
@@
The contribution of the $O$-terms is $O(X^{3/2+\varepsilon})$ by Prop. 5.4. We now analyze the exponential integral
@@
Let us assume further $m,n>0$ since the analysis in the other case is similar and we get the same bound. We consider two cases.
Case I: $x_{0}\in[2,2la_{1}/3]$. Here we split the interval $1,la_{1}]$ in three parts as follows:
@@ 7.1. Contributions of $I_{4}$
By Prop. 5.4, the contributions of the $O$-terms are absorbed into the error term. Define
$I_{m,n,a_{1}}=\int_{la_{1}}^{X}e(h(x))\,\textnormal{d}x,\quad\textnormal{ where }\quad h(x)=-\frac{nx}{a_{1}}-\frac{mX}{x}.$
@@
If $mn<0$, then there is no stationary point we easily get the bound $I_{m,n,a_{1}}\ll\frac{a_{1}}{n}$ and the contribution of this term is sufficiently small by Prop. 5.4. The case left to consider is $mn>0$ and we shall assume that both $m$ and $n$ are positive since the complimentary case will be similar. In this case the stationary point is given by
@@
Now, if we give an upper bound for exponential integrals and then sum over the three variables $m,n,$ and $a_{1}$ as in the case of $I_{1}$ then the error term we get is not good enough. We, therefore, refine our analysis by localizing the stationary point $x_{0}$ and extracting the stationary phases and estimating the ensuing error terms carefully. This leads to the following:
@@
Before we proceed with the proof, let us first observe the following crude range for $x_{0}$: $x_{0}\ll X\sqrt{X/H}\ll X^{5/4}$. This is immediate from (7.2).
@@ 7.2. Localizing the stationary point and decomposition of $B_{1,w}(X,I_{4})$
For a fixed $a_{1}$, define
$k=k(a_{1})$ $=\max\left\{j\geq 0:\frac{X}{2^{j}}\geq la_{1}+1\right\}=\left[\frac{\log(X/(la_{1}+1))}{\log 2}\right].$ (7.3)
@@
$\text{for }0\leq j\leq k=k(a_{1})\leq J(X),\text{ we have }X/2^{j}\geq la_{1},\text{ where }J(X):=\left[\frac{\log(X/2)}{\log 2}\right].$ (7.4)
We make a dyadic subdivision of the range of $x_{0}$, i.e., the interval $1,X^{5/4}]$ in the following manner:
$\mathcal{I}_{0}=\left[X+\delta,X^{\frac{5}{4}}\right],\mathcal{I}_{j}=\left[\frac{X}{2^{j}}+\delta,\frac{X}{2^{j-1}}-\delta\right]\text{ for }1\leq j\leq k,\text{ and }\mathcal{I}_{k+1}=[la_{1}+\delta,\frac{X}{2^{k}}-\delta],\mathcal{I}_{k+2}=[1,la_{1}-\delta].$ (7.5)
@@
Here $\delta>0$ is to be chosen later (we will see that $\delta=1/100$ suffices). Clearly, we can write the interval $1,X^{5/4}]$ as a disjoint union of the $\mathcal{I}_{j}$’s and $\mathcal{I}^{\prime}_{j}$s, and thus we decompose the sum as follows:
$B_{1,w}(X,I_{4})$ $=\frac{1}{2\pi i}\sum_{l|r}\sum_{a_{1}>0}\sum_{0\leq j\leq k+2}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}\frac{1}{m}\frac{w(la_{1})S(nr_{1},-m,a_{1})}{a_{1}}I_{m,n,a_{1}}$ (7.7)
@@
where in the sum $\mathcal{S}$, the stationary point $x_{0}\in\cup_{j}\mathcal{I}_{j}$ and in the sum $\mathcal{T}$, $x_{0}\in\cup_{j}\mathcal{I}^{\prime}_{j}$.
7.3. Subdivision of the range of integration and decomposition of the sum $\mathcal{S}$
Let
$\mathcal{F}_{j}=\left[\frac{X}{2^{j}},\frac{X}{2^{j-1}}\right],\text{ for }1\leq j\leq k\text{ and }\mathcal{F}_{k+1}:;=\left[la_{1},\frac{X}{2^{k}}\right].$ (7.8)
Clearly, $la_{1},X]=\bigcup\limits_{1\leq j\leq k+1}\mathcal{F}_{j},$ where an interval $\mathcal{F}_{j}$ can intersect its neighbour only at the boundary point. Let us define